OpenStudy (anonymous):

What is the slope of the line y - 6 = 2(x + 4).

OpenStudy (anonymous):

y = mx + x m = slope of the libe

OpenStudy (gorv):

y=mx+c standard form

OpenStudy (gorv):

coefficient of x = slope

OpenStudy (gorv):

what is coefficient of x in your equation??

OpenStudy (imtiaz7):

y-6=2(x+4) y-6=2x+8 y=2x+8+6 y=2x+14 we know that slope intercept form is y=mx+b compare these equation then you will get the slope that is always with x ie m

OpenStudy (imtiaz7):

i think there is no need of further explanation to find the slope?

OpenStudy (imtiaz7):

Good!